OpenStudy (anonymous):

Can 5 9/6 be smiplified

OpenStudy (anonymous):

Is that \[5\frac{ 9 }{ 6 }\]

OpenStudy (anonymous):

Yes

OpenStudy (jhannybean):

Make the complex fraction into a simpple fraction by: \(\dfrac{(6\cdot 5) +9}{6}\)

OpenStudy (anonymous):

6 divided by 39 i mean

OpenStudy (anonymous):

39 / 6

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Now convert it back to a mixed fraction, how many 6 fit in 39

OpenStudy (anonymous):

6

OpenStudy (anonymous):

So that would be your wholes, now what are you left over with?

OpenStudy (anonymous):

5 and 6/6?

OpenStudy (anonymous):

39 - 36 = 3

OpenStudy (anonymous):

5 3/6

OpenStudy (anonymous):

6 fits in 39, 6 times, with a remainder of 3 divided by 6

OpenStudy (anonymous):

and that's 2

OpenStudy (anonymous):

\[6\frac{ 3 }{ 6 }\]

OpenStudy (anonymous):

or simplified even further: \[6\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I got it now.